# Noise-Aware Entanglement Generation Protocols for Superconducting Qubits with impedance-matched FBAR Transducers

### v2

@@ II.3 Modifying the equivalent circuit to incorporate a photonic molecule
Here we modify the extended BVD model to include the optomechanical loading from a photonic molecule, i.e. a system of two coupled micro-ring resonators, for the first time. A photonic molecule is a two-level photonic system that can be dynamically controlled by gigahertz-frequency microwave signals, and it typically consists of two coupled MRRs, where only one MMR is coupled to a bus waveguide [58]. This extension is a crucial step for modeling a wide range of optomechanical transducers, since many rely on the mode splitting within a photonic molecule to achieve triply resonant transducer operation [38]. When two MRRs are coupled at a rate $J$, their modes hybridize into supermodes. So-called symmetric and anti-symmetric supermodes appear with a frequency splitting equal to $2J$. For microwave-to-optical transduction, the optical pump is tuned to the lower-frequency symmetric supermode frequency $\omega_{s}$, which sits $\omega_{m}=2J$ lower than $\omega_{A}$. This triply-resonant condition, when satisfied, should boost the conversion efficiency at a given optical power due to the cavity enhancement of the pump tone [38].
@@
$R_{OM}^{eq}=\left(\frac{1}{R_{OM}}+\frac{1}{R_{OO,+}-R_{OO,-}}\right)^{-1}\\ =R_{OM,+}^{eq}-R_{OM,-}^{eq}=\frac{R_{OM}(R_{OO,+}-R_{OO,-})}{R_{OM}+R_{OO,+}-R_{OO,-}}\\ =R_{m}\mathcal{C}_{OM}^{eq}(\mathscr{L}_{+}^{2}-\mathscr{L}_{-}^{2})$ (41)
@@
What expressions should we use to extract the device’s figures of merit now? The expression for thermal noise remains unchanged (see Equation 29), but the expressions for Raman noise, conversion efficiency and bandwidth do change. The optical Raman noise is now
@@
II.4 Matching circuits for single-ring and double-ring FBAR transducers
@@
II.4.1 Matching Circuit for Single-Ring FBAR Transducer
@@
When designing impedance matching networks, especially those that require $C_{T}=0$ to minimize added noise, we must take care to consider the effects of parasitic capacitance terms. Spiral inductors, like the one depicted in Ref. [27], have significant self-capacitances and capacitances to ground. If ignored, stray capacitances will pull the matching circuit away from its optimal regime, leading to degraded device performance. Likewise, matching capacitors will have parasitic inductances. In Appendix C we investigate the effects of these unwanted terms on the matching network and transducer performance.
II.4.2 Matching Circuit for Double-Ring FBAR Transducer
@@
II.4.3 Example Matching Circuit Design
@@
The example designs are shown with a $200\;\mathrm{nm}$ thick Niobium superconducting layer, which defines the superconducting circuit and the ground plane. A wire-bonding pad (top of panel (b)) connects to a $50\;\mathrm{Ohm}$ superconducting waveguide, which connects galvanically to the top electrode of the transducer. This electrode is also galvanically connected to the impedance matching circuit to achieve the in-parallel configuration dictated by the extended BVD circuit model. The ground plane for the impedance matching circuit is the same as the ground of the transducer’s piezo cavity. The efficiency-maximizing version of the impedance matching network shown in panel (a,b) incorporates a Niobium meandering inductor. The total footprint is an area of approximately $100\;\mathrm{\mu m}\times 100\;\mathrm{\mu m}$. The meander line width, center-to-separation, and total length should be designed to target a total inductance in the $2-4\;\mathrm{nH}$ range with an anticipated parasitic capacitance in the $5-10\;\mathrm{fF}$ range. The specific lumped element values will depend on the quality of the deposited materials and fabrication conditions, so this example is meant to be qualitiative and not exact. The matching capacitor is designed in a rectangular shape with dimensions of approximately $300\;\mathrm{\mu m}\times 325\;\mathrm{\mu m}$, and has an additional capacitance of $400-500\;\mathrm{fF}$.
A device with a noise-minimizing version of the matching network is shown in panel (c). Since this matching network requires a capacitance as close to zero as possible, the square capacitor is discarded and the meandering inductor is swapped out for a high kinetic inductance nanowire inductor [61]. The nanowire inductor shown consists of a $\sim 20\;\mathrm{nm}$-thick NbN layer with a width of $1\;\mathrm{\mu m}$ and length of $\sim 150\;\mathrm{\mu m}$. Depending on the specific materials and fabrication recipes involved, the nanowire length can possibly be reduced to $<10\;\mathrm{\mu m}$ to achieve an inductance in the $2-4\;\mathrm{nH}$ range with an estimated parasitic capacitance is in the $\mathrm{aF}$ range. Additional discussion exploring the effects of parasitic lumped element terms can be found in Appendix C.
@@ III BVD-Informed Entanglement Heralding Protocols
Now that we have established how to design impedance matching networks for the FBAR transducer, both to maximize conversion efficiency and minimize added noise, we investigate how the impedance-matched devices perform within Type-I and Type-II entanglement heralding protocols. Readout fidelities and entanglement generation rates are plotted and analyzed in various experimental operating conditions. These results allow us to make informed decisions in the device design phase to optimize the chosen quantum networking protocol, given a set of realistic experimental parameters such as temperature, transducer pump power, and link loss between the transducer and the Bell state analyzer.
Let us begin with Type-I entanglement heralding. Because this is a single-photon protocol, it will succeed only when one photon is emitted, in total, from both network nodes (see Fig. 2). For each trial of the experiment, there is a small chance that both superconducting network nodes will produce an entangled telecom photon, and that one of these photons will be lost in the network before reaching the photodetector. Instances such as these, where two photons are emitted but only one is detected, and the two events are completely uncorrelated, produce false positive events and contribute to the infidelity. In conceiving false positive detection events this way, we are assuming that two-photon detection events can be sorted from one-photon detection events and discarded through post-selection.
While this source of infidelity is limiting, in practice, the primary downside of Type-I entanglement is the requirement of absolute relative phase stability between the photon sources and the beamsplitter [48]. Slight changes in optical fiber path length due to temperature fluctuations, for example, can change the relative phase of the entangled state. This approach could still be suitable for entanglement generation between cryostats in the same building, e.g. local area networks, but will quickly become untenable for longer networks with larger inherent phase instabilities. In the discussion below, the fidelity, or readout fidelity, of the protocol refers to the conditional fidelity of Bell state generation [28], that is, the overlap of the created quantum state with the desired quantum state, given that the relevant photodetector click condition is fulfilled.
@@
Assuming a Poissonian photon detection process, following the Stochastic master equation [12], a photon is emitted by the transducer with probability
$P_{1}=1-P_{0}=1-e^{-r_{o}\Delta t}$ (48)
where $r_{o}$ is the photon generation rate and $\Delta t$ is the fixed time duration of the transduction process. Note that Eq. 48 represents the probability that at least one photon is emitted, and not precisely one photon. The probability of precisely one photon emission event occurring within $\Delta t$ is
@@
Next we define the expression for $\Delta t$. For an FBAR transducer with a single optical cavity, we have the transduction time duration
@@
which can be understood as the sum of the lifetimes of each of the constituent cavity modes of the device, plus the time it takes for a transducer photon to couple out of the MRR and into the optical bus waveguide. For a two-MRR transducer we have
@@
|$\ket{\Psi_{SQ}}=\sqrt{P_{01}}\ket{e_{A}g_{B}}+\sqrt{P_{10}}\ket{g_{A}e_{B}}$ (56)
where |$\ket{g}\;,\;\ket{e}$ are the qubits’ ground and excited states, respectively. During an entanglement generation attempt, there is a probability that both network nodes will emit entangled photons, but one photon is lost before reaching the photodetector. This scenario contributes false positive events to the infidelity of the protocol, since both superconducting qubits will be in their ground states, even though we think one is in its excited state. This occurs with $P_{in}$, the probability of an infidelity-causing event, of approximately
@@
As such, the fidelity of the protocol is then:
@@
In other words, the fidelity is equal to the ratio of the total probability of creating the desired state to the sum of the probabilities of creating the desired state and of creating an undesired state with a single detection event. As the pump power to the transducers increases, $P_{1}^{2}$ increases, and the fidelity is degraded. Note that by counting false positive events in this manner, we assume that we are able to distinguish between single count events (desired) and coincident events (undesired) and discard those through post-selection. The associated photon generation rate is
@@ III.1 Incorporating Infidelity Contributions Due to Thermal Noise
Here is the point where the extended BVD model of our transducer comes into play. For a given set of device and operating parameters, the BVD model outputs matching circuit parameters, an estimated conversion efficiency, and estimates for added noise quanta due to Raman and thermal noise. (Note that in this section we disregard Raman noise quanta, since there are multiple orders of magnitude fewer noise quanta due to Raman noise.) We can use the thermal noise output (Equation 29) from the BVD model as an input to our entanglement generation fidelity expressions, while simultaneously optimizing our transducer for either highest-efficiency or lowest-noise operation. To do this, we input the BVD value of $n_{th}$ into Equation 62.
@@ III.2 Two-Photon Protocol with Thermal Noise
Table 2: Example results for FBAR transducers (for $n_{cav}=625\;\mathrm{mil}$ or $g_{OM}=10\;\mathrm{MHz}$ and $\Delta=\omega_{s}$ at $50\;\mathrm{mK}$).
@@
[…]
@@
Motional resistance $R_{m}$ $44.588\;\mathrm{\Omega}$ $44.588\;\mathrm{\Omega}$ $44.588\;\mathrm{\Omega}$ $44.588\;\mathrm{\Omega}$
Lorentzian upper sideband $\mathscr{L}_{+}^{2}$ $1$ $1$ $1$ $1$
Lorentzian lower sideband $\mathscr{L}_{-}^{2}$ $41.303E-$ 4 $41.303E-$ 4 $41.303E-$ 4 $41.303E-$ 4
@@
Lower sideband resistance 1 ring $R_{OM,-}$ $5.959E-3\;\mathrm{\Omega}$ $5.959E-3\;\mathrm{\Omega}$ N/A N/A
@@
Upper sideband resistance 2 ring $R_{OO,+}$ N/A N/A $22.908\;\mathrm{k\Omega}$ $22.908\;\mathrm{k\Omega}$
@@
[…]
@@
Raman noise quanta $n_{o}$ $56.598E-$ 5 $51.829E-$ 5 $56.591E-$ 5 $51.827E-$ 5
Thermal noise quanta $n_{th}$ $22.203E-$ 2 $36.110E-$ 3 $22.205E-$ 2 $36.112E-$ 3
Figure 9: Entanglement readout fidelity for Type-I and Type-II versus link efficiency for (a)$g_{OM}=1\;\mathrm{MHz}$ and (b) $g_{OM}=10\;\mathrm{MHz}$ at $T=50\;\mathrm{mK}$. Red curves: Type-I, minimize noise, Blue curves: Type-I, maximize efficiency Gold curves: Type-II, minimize noise, Green curves: Type-I, maximize efficiency.
Table 3: Example results for a single-ring FBAR transducer, $T=50\;\mathrm{mK}$, $g_{OM}=10\;\mathrm{MHz}$ ($625\times 10^{6}$ photons in optical cavity), $\eta_{l}=50\%$. Qubit reset time is $1\;\mathrm{\mu s}$, time-bin separation $t_{sep}=\Delta_{t}+t_{r}$ and detector efficiency is $90\%$.
@@
[…]
Photon emission probability $P_{1}$ $24.364\%$ $22.947\%$ $24.364\%$ $22.947\%$
@@
[…]
@@
Table 4: Example results for a double-ring FBAR transducer, $T=50\;\mathrm{mK}$, $g_{OM}=10\;\mathrm{MHz}$ ($625\times 10^{6}$ photons in optical cavity), $\eta_{l}=50\%$. Qubit reset time is $1\;\mathrm{\mu s}$, time-bin separation $t_{sep}=\Delta_{t}+t_{r}$ and detector efficiency is $90\%$.
@@
[…]
@@
[…]
Fidelity F $90.640\%$ $92.504\%$ $92.817\%$ $97.899\%$
Figure 10: Type I and Type-II Entanglement fidelity and rate for maximum-efficiency and minimum-noise matching networks, versus temperature, for (a)$g_{OM}=1\;\mathrm{MHz}$ and $\eta_{l}=0.5$, (b)$g_{OM}=1\;\mathrm{MHz}$ and $\eta_{l}=0.1$, (c) $g_{OM}=10\;\mathrm{MHz}$ and $\eta_{l}=0.5$ and (d) $g_{OM}=10\;\mathrm{MHz}$ and $\eta_{l}=0.1$. The entanglement fidelity degrades for the Type-I protocol at higher pump powers (higher $g_{OM}$), though using a noise-minimizing matching circuit mitigates this effect somewhat. At high temperatures, the fidelity is the worst for the efficiency-maximizing devices, regardless of whether a Type-I or Type-II protocol is used. At low temperatures, the optimal configuration depends on $g_{OM}$. For a realistic link efficiency of $\eta_{l}=0.1$ (b,d), the Type-II protocol with a noise-minimizing matching circuit results in the highest entanglement fidelity at $<50\;\mathrm{mK}$. Plots (e,f) Show the entanglement generation rate versus $g_{OM}$ for (e) $\eta_{l}=0.5$ link loss and (f)$\eta_{l}=0.1$. As anticipated, Type-II protocols achieve lower entanglement rates than Type-I protocols.
@@
For the Type-II infidelity, do we also need to include terms corresponding to thermal photon emission in addition to the emission of two signal photons, e.g. a three-photon emission event, where one of the emitted photons is lost before reaching the photodetectors? Fortunately, the answer is no. If we have a situation where the two-click condition is satisfied by one signal photon each in the Early and Late time bins, and we have successfully generated the distributed superconducting qubit Bell pair (Equation 67), the presence of an additional thermal phonon within the transducer will not affect the distributed Bell state. The emissive nature of the direct transduction protocol protects us from this potential source of readout infidelity, assuming that the superconducting qubit is protected by the spurious phonon excitation via a tunable coupling element [66] placed between the transducer and the superconducting qubit that protects the qubit from spurious microwave photon absorption. This protection is not necessarily present for a blue-detuned protocol, as discussed in Appendix E.

### v3

@@ II.3 Modifying the equivalent circuit to incorporate a photonic molecule
Here we modify the extended BVD model to include the optomechanical loading from a photonic molecule, i.e. a system of two coupled micro-ring resonators, for the first time. A photonic molecule is a two-level photonic system that can be dynamically controlled by gigahertz-frequency microwave signals, and it typically consists of two coupled MRRs, where only one MMR is coupled to a bus waveguide [59]. This extension is a crucial step for modeling a wide range of optomechanical transducers. While single-ring FBAR transducers have demonstrated promising performance for transduction, many leading transducers incorporate double-ring optical resonators. Double-ring resonators enable more straightforward bidirectional operation[38] and, in some cases, frequency tunability of the hybridized optical supermodes [60, 61]. Many rely on the tunable mode splitting within a photonic molecule to achieve triply resonant transducer operation [38]. Beyond this, the supermode tunability can allow for frequency matching of the emission from two separate transducers, at least on small frequency scales, to correct for modest fabrication variations across separate devices that may impact the degree of in-distinguishability required for interference. Since we aim to incorporate these transducers into interference-based quantum networking protocols, there is an urgent need to generalize the impedance matching formalism to the double-ring class of devices. See Appendices I, H for in-depth discussions of these issues.
When two MRRs are coupled at a rate $J$, their modes hybridize into supermodes. So-called symmetric and anti-symmetric supermodes appear with a frequency splitting equal to $2J$. For microwave-to-optical transduction, the optical pump is tuned to the lower-frequency symmetric supermode frequency $\omega_{s}$, which sits $\omega_{m}=2J$ lower than $\omega_{A}$. This triply-resonant condition, when satisfied, should boost the conversion efficiency at a given optical power due to the cavity enhancement of the pump tone [38].
@@
$R_{OM}^{eq}=\left(\frac{1}{R_{OM}}+\frac{1}{R_{OO,+}-R_{OO,-}}\right)^{-1}\\ =R_{OM,+}^{eq}-R_{OM,-}^{eq}\\ =R_{m}\mathcal{C}_{OM}^{eq}(\mathscr{L}_{+}^{2}-\mathscr{L}_{-}^{2})$ (41)
@@
What expressions should we use to extract the device’s figures of merit now? The expression for thermal noise remains unchanged (see Equation 29), but the expressions for Raman noise, conversion efficiency and bandwidth do change. The optical Raman noise $n_{o}$ is now
@@
II.4 Incorporating parasitic circuit elements
The formalism developed so far, like previous work, ignores stray capacitances associated with the matching inductor $L$ in the equivalent circuit model. This is despite the large matching inductances and small matching capacitances often required for OMC-based transducers [27], which may require large spiral inductors that have large stray, or parasitic, capacitances. The stray inductance of the matching capacitor should also be considered. As explored in-depth in Appendix D, we derive updated expressions for the matching capacitance and matching inductance that account for parasitic terms. The modified matching terms ($C_{T}$’, $L$’) can be written as:
$C_{T}^{\prime}=C_{T}+C_{p}$ (48)
and
$L^{\prime}=L+L_{CT}.$ (49)
An updated equivalent circuit diagram incorporating elements for parasitic terms is shown in Fig.  6.
Figure 6: A modified equivalent circuit for a transducer with parasitic terms incorporated into the matching network. $L_{CT}$: parasitic inductance of the matching capacitor. $C_{s}$: parasitic capacitance of the matching inductor.
II.5 Matching circuits for single-ring and double-ring FBAR transducers
@@
II.5.1 Matching Circuit for Single-Ring FBAR Transducer
@@
When designing impedance matching networks, especially those that require $C_{T}=0$ to minimize added noise, we must take care to consider the effects of parasitic capacitance terms. Spiral inductors, like the one depicted in Ref. [27], have significant self-capacitances and capacitances to ground. If ignored, stray capacitances will pull the matching circuit away from its optimal regime, leading to degraded device performance. Likewise, matching capacitors will have parasitic inductances. In Appendix D we investigate the effects of these unwanted terms on the matching network and transducer performance.
II.5.2 Matching Circuit for Double-Ring FBAR Transducer
@@
II.5.3 Example Matching Circuit Design
@@
The example designs are shown with a $200\;\mathrm{nm}$ thick Niobium superconducting layer, which defines the superconducting circuit and the ground plane. A wire-bonding pad (top of panel (b)) connects to a $50\;\mathrm{Ohm}$ superconducting waveguide, which connects galvanically to the top electrode of the transducer. This electrode is also galvanically connected to the impedance matching circuit to achieve the in-parallel configuration dictated by the extended BVD circuit model. The ground plane for the impedance matching circuit is the same as the ground of the transducer’s piezo cavity. The efficiency-maximizing version of the impedance matching network shown in panel (a,b) incorporates a Niobium meandering inductor. The total footprint is an area of approximately $100\;\mathrm{\mu m}\times 100\;\mathrm{\mu m}$. The meander line width, center-to-separation, and total length should be designed to target a total inductance in the $2-4\;\mathrm{nH}$ range with an anticipated parasitic capacitance in the $5-10\;\mathrm{fF}$ range. The specific lumped element values will depend on the quality of the deposited materials and fabrication conditions, so this example is meant to be qualitative and not exact. The matching capacitor is designed in a rectangular shape with dimensions of approximately $300\;\mathrm{\mu m}\times 325\;\mathrm{\mu m}$, and has an additional capacitance of $400-500\;\mathrm{fF}$.
A device with a noise-minimizing version of the matching network is shown in panel (c). Since this matching network requires a capacitance as close to zero as possible, the square capacitor is discarded and the meandering inductor is swapped out for a high kinetic inductance nanowire inductor [64]. The nanowire inductor shown consists of a $\sim 20\;\mathrm{nm}$-thick NbN layer with a width of $1\;\mathrm{\mu m}$ and length of $\sim 150\;\mathrm{\mu m}$. Depending on the specific materials and fabrication recipes involved, the nanowire length can possibly be reduced to $10\;\mathrm{\mu m}$ to achieve an inductance in the $2-4\;\mathrm{nH}$ range with an estimated parasitic capacitance is in the $\mathrm{aF}$ range. Additional discussion exploring the effects of parasitic lumped element terms can be found in Appendix D. The impacts of fabrication variances of the matching circuit elements on transducer and entanglement heralding performance are discussed in Appendix J. A sensitivity analysis reveals that the FBAR transducers are robust to realistic fabrication tolerances, as are the quantum networking protocols that rely on these devices.
@@ III BVD-Informed Entanglement Heralding Protocols
Now that we have established how to design impedance matching networks for the FBAR transducer, both to maximize conversion efficiency and minimize added noise, we investigate how the impedance-matched devices perform within Type-I and Type-II entanglement heralding protocols. The generalized extended BVD model developed in Section II provides guidance for our current task by connecting device-level optimization to protocol-level performance. The above formalism enables systematic selection of optimal transducer operating points that balance conversion efficiency and added noise for maximum entanglement fidelity. As we will show, this framework allows quantum network engineers to make informed trade-offs between entanglement generation rate and fidelity based on application requirements.
Toward this end, conditional readout fidelities and entanglement generation rates are plotted and analyzed in various experimental operating conditions. These results allow us to make informed decisions in the device design phase to optimize the chosen quantum networking protocol, given a set of realistic experimental parameters such as temperature, transducer pump power, and link loss between the transducer and the Bell state analyzer.
Here we define the conditional readout fidelity $F$ as the overlap between the generated quantum state and the target Bell state, given that the proper heralding click condition is satisfied:
[…]
Let us begin with Type-I entanglement heralding. Because this is a single-photon protocol, it will succeed only when one photon is emitted, in total, from both network nodes (see Fig. 2). For each trial of the experiment, there is a small chance that both superconducting network nodes will produce an entangled telecom photon, and that one of these photons will be lost in the network before reaching the photodetector. Instances such as these, where two photons are emitted but only one is detected, and the two events are completely uncorrelated, produce false positive events and contribute to the infidelity (i.e., the decrease in fidelity). In classifying false positive detection events this way, we are assuming that two-photon detection events can be sorted from one-photon detection events and discarded through post-selection.
While this source of infidelity is limiting, in practice, the primary downside of Type-I entanglement heralding is the requirement of absolute relative phase stability between the photon sources and the beamsplitter [49]. Slight changes in optical fiber path length due to temperature fluctuations, for example, can change the relative phase of the entangled state. This approach could still be suitable for entanglement generation between cryostats in the same building, e.g. local area networks, but will quickly become untenable for longer networks with larger inherent phase instabilities. In the discussion below, the fidelity, or readout fidelity, of the protocol refers to the conditional fidelity of Bell state generation [28], that is, the overlap of the created quantum state with the desired quantum state, given that the relevant photodetector click condition is fulfilled.
@@
Assuming a Poissonian photon detection process, following the Stochastic master equation [12], a photon is emitted by the transducer with approximate probability
$P_{1}\approx 1-P_{0}=1-e^{-r_{o}\Delta t}$ (51)
where $r_{o}$ is the photon generation rate and $\Delta t$ is the characteristic time duration of the transduction process. Note that Eq. 51 represents the probability that at least one photon is emitted, and not precisely one photon. The probability of precisely one photon emission event occurring within $\Delta t$ is
@@
Next we define the expression for $\Delta t$. For an FBAR transducer with a single optical cavity, we have the characteristic transduction time duration
@@
which can be understood as the sum of the lifetimes of each of the constituent cavity modes of the device, including the time it takes for a transducer photon to couple out of the MRR and into the optical bus waveguide. For a two-MRR transducer we have
@@
[…]
@@
As such, the fidelity of the protocol following Equation 50 is then:
@@
In other words, the fidelity is equal to the ratio of the total probability of creating the desired state to the sum of the probabilities of creating the desired state and of creating an undesired state with a single detection event. As the pump power to the transducers increases, $P_{1}^{2}$ increases, and the fidelity is degraded. Note that by counting false positive events in this manner, we assume that we are able to distinguish between single count events (desired) and coincident events (undesired) and discard those through post-selection. The associated photon generation rate $n_{r}$ is
@@
This analysis assumes that the achievable readout fidelity is not reduced by spatio-temporal mode mismatch between the photons emitted by the two nodes. As explored in Appendix H, within expected fabrication tolerances, we anticipate the spatio-temporal overlap from photons emitted by nominally identical transducers to achieve fidelity ceilings of $99\%.$
@@ III.1 Incorporating Infidelity Contributions Due to Thermal Noise
For the interested reader, a detailed breakdown of Equations 66, 67 is available in Appendix F. Here is the point where the extended BVD model of our transducer comes into play. For a given set of device and operating parameters, the BVD model outputs matching circuit parameters, an estimated conversion efficiency, and estimates for added noise quanta due to Raman and thermal noise. (Note that in this section we disregard Raman noise quanta, since there are multiple orders of magnitude fewer noise quanta due to Raman noise.) We can use the thermal noise output (Equation 29) from the BVD model as an input to our entanglement generation fidelity expressions, while simultaneously optimizing our transducer for either highest-efficiency or lowest-noise operation. To do this, we input the BVD value of $n_{th}$ into Equation 65.
@@ III.2 Two-Photon Protocol with Thermal Noise
Table 2: Example results for FBAR transducers (for $n_{cav}=6.25\times 10^{8}$ or $g_{OM}=10\;\mathrm{MHz}$ and $\Delta=\omega_{s}$ at $50\;\mathrm{mK}$).
@@
Motional capacitance $C_{m}$ $0.860\;\textrm{fF}$ $0.860\;\textrm{fF}$ $0.860\;\textrm{fF}$ $0.860\;\textrm{fF}$
@@
Motional resistance $R_{m}$ $4.588\;\mathrm{\Omega} 44$$588\;\mathrm{\Omega} 44.$$588\;\mathrm{\Omega} 44.$$4.588\;\mathrm{\Omega}$
Lorentzian upper sideband $\mathscr{L}_{+}^{2}$ 11 11 11 11
Lorentzian lower sideband $\mathscr{L}_{-}^{2}$ $3\times 10^{-4} 1.303×10$$3\times 10^{-4}$ $3\times 10^{-4} 1.303×10$$3\times 10^{-4}$
@@
Lower sideband resistance 1 ring $R_{OM,-}$ $59\times 10^{-3}\;\mathrm{\Omega} 5.959$$59\times 10^{-3}\;\mathrm{\Omega}$ N/A N/A
@@
Upper sideband resistance 2 ring $R_{OO,+}$ N/A N/A $2.908\;\mathrm{k\Omega} 22$$2.908\;\mathrm{k\Omega}$
@@
Matching capacitance $C_{T}$ $522.346\;\textrm{fF}$ $0\;\textrm{fF}$ $522.680\;\textrm{fF}$ $0\;\textrm{fF}$
@@
Raman noise quanta $n_{o}$ $6.598\times 10^{-5}$ $1.829\times 10^{-5}$ $6.591\times 10^{-5}$ $1.827\times 10^{-5}$
Thermal noise quanta $n_{th}$ $2.203\times 10^{-2}$ $6.110\times 10^{-3}$ $2.205\times 10^{-2}$ $6.112\times 10^{-2}$
[…]
Table 3: Example results for a single-ring FBAR transducer, $T=50\;\mathrm{mK}$, $g_{OM}=10\;\mathrm{MHz}$ ($6.25\times 10^{8}$ photons in optical cavity), $\eta_{l}=50\%$. Qubit reset time is $1\;\mathrm{\mu s}$, time-bin separation $t_{sep}=\Delta_{t}+t_{r}$ and detector efficiency is $90\%$.
@@
Transduction time $\Delta t$ $25.655\;\textrm{ns} 11$$17.304\;\textrm{ns}$ $25.655\;\textrm{ns} 11$$17.304\;\textrm{ns}$
Photon emission rate $r_{o}$ $22\;\textrm{MHz} 2.222$$\textrm{MHz} 2.222$$\textrm{MHz} 2.222$$22\;\textrm{MHz}$
Photon emission probability $P_{1}$ $4.364\% 22$$2.947\%$ $4.364\% 22$$2.947\%$
@@
Photon generation rate $n_{r}$ or $n_{r}^{\prime}$ $375.252\;\textrm{kHz}$ $359.542\;\textrm{kHz}$ $187.626\;\textrm{kHz}$ $179.771\;\textrm{kHz}$
Entanglement rate $\tau_{ent}$ $168.863\;\textrm{kHz}$ $161.794\;\textrm{kHz}$ $37.994\;\textrm{kHz}$ $36.403\;\textrm{kHz}$
@@
Table 4: Example results for a double-ring FBAR transducer, $T=50\;\mathrm{mK}$, $g_{OM}=10\;\mathrm{MHz}$ ($6.25\times 10^{8}$ photons in optical cavity), $\eta_{l}=50\%$. Qubit reset time is $1\;\mathrm{\mu s}$, time-bin separation $t_{sep}=\Delta_{t}+t_{r}$ and detector efficiency is $90\%$.
@@
Transduction time $\Delta t_{2MRR}$ $26.248\;\textrm{ns} 11$$17.893\;\textrm{ns}$ $26.248\;\textrm{ns} 11$$17.893\;\textrm{ns}$
Photon emission rate $r_{o}$ $22\;\textrm{MHz} 2.222$$\textrm{MHz} 2.222$$\textrm{MHz} 2.222$$22\;\textrm{MHz}$
@@
Photon generation rate n $376.327\;\textrm{kHz}$ $360.683\;\textrm{kHz}$ $188.163\;\textrm{kHz}$ $180.341\;\textrm{kHz}$
Entanglement rate $\tau_{ent}$ $169.347\;\textrm{kHz}$ $162.307\;\textrm{kHz}$ $38.103\;\textrm{kHz}$ $36.519\;\textrm{kHz}$
Fidelity F $90.640\%$ $92.504\%$ $92.817\%$ $97.899\%$
@@
For the Type-II infidelity, do we also need to include terms corresponding to thermal photon emission in addition to the emission of two signal photons, e.g. a three-photon emission event, where one of the emitted photons is lost before reaching the photodetectors? Fortunately, the answer is no. If we have a situation where the two-click condition is satisfied by one signal photon each in the Early and Late time bins, and we have successfully generated the distributed superconducting qubit Bell pair (Equation 70), the presence of an additional thermal phonon within the transducer will not affect the distributed Bell state. The emissive nature of the direct transduction protocol protects us from this potential source of readout infidelity, assuming that the superconducting qubit is protected by the spurious phonon excitation via a tunable coupling element [69] placed between the transducer and the superconducting qubit that protects the qubit from spurious microwave photon absorption. This protection is not necessarily present for a blue-detuned protocol, as discussed in Appendix G.